OpenStudy (anonymous):

eigenvector help needed I have the eigenvalues(-1,1,2,4) but I need help with the eigenvector for 2 I Will post the matrix

OpenStudy (anonymous):

post the matrix!

OpenStudy (anonymous):

\[\left[\begin{matrix}1 & 0 & 0&0 \\ 1&2 &0&0 \\ 1&2&-1&1 \\ 1&1&0&4\end{matrix}\right]\]

OpenStudy (anonymous):

sorry that took so long

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Whn I put it into wolfram alpha, it says I should get (0,-2,-1,1) but I can't figure out how to get that. I got all the other ones, but this one isn't coming out.

OpenStudy (amistre64):

so that thing you posted is the matrix you are working with right?

OpenStudy (anonymous):

yep

OpenStudy (amistre64):

and you got the values of: 4,2,-1,1 correct?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

I just can't figure out how to get the correct eigenvector for 2. I got the other ones but this one isn't coming out correctly

OpenStudy (anonymous):

the answer (i made it from my hand) is (0, 2, 1, -1) and the wólfram alpha is right. Simply find A-2I where I is the identity matrix. Row-reduce it (applying elementary operations in the rows). You get (1,0,0,0//0,1,0,2//0,0,1,1//0,0,0,0)=B. Solving BX=O produces the eigenvector.

OpenStudy (anonymous):

wait how did you get B? I'm not quite seeing where you got that from

OpenStudy (anonymous):

Row-reducing A-2I

OpenStudy (anonymous):

it may just be a different name but what do you mean by row reducing?

OpenStudy (anonymous):

apply one of the 3 elementary operations in rows, to obtain the ECHELON form of the matrix

OpenStudy (amistre64):

(A-L)x = 0, should produce the eigenvectors for x, right?

OpenStudy (anonymous):

1. Interchange rows 2. Multiply a row by a no zero number 3. Add a row, multiplied by a number, to another row

OpenStudy (anonymous):

amistre64 is right. It's A-I, not A-L (what is L???)

OpenStudy (amistre64):

L is lamda :) a given eigenvalue

OpenStudy (anonymous):

right. a-2I gets (-1,0,0,0//1,0,0,0//1,0,-3,1//1,1,0,0), right?

OpenStudy (anonymous):

Noooo A-L means a matrix (A) MINUS a number (L), you can't do this

OpenStudy (anonymous):

lambda times the identity matrix, yes?

OpenStudy (amistre64):

im assuming that since we are talking about eigen stuff, that the I is implied ...

OpenStudy (anonymous):

better: (A-LI)x=o, where I = identity matrix

OpenStudy (anonymous):

a- 2 (lambda) times the identity matrix

OpenStudy (anonymous):

please do not "imply".

OpenStudy (anonymous):

excuse my english

OpenStudy (anonymous):

anyway... what I put earlier is what you would get, right?

OpenStudy (anonymous):

goalie is right

OpenStudy (anonymous):

so then why can I not seem to get the correct vector from that matrix?

OpenStudy (anonymous):

i have to go... nice to meet you... Greetings from La Paz, Bolivia!!

OpenStudy (anonymous):

... I still haven't figured out why I can't seem to get the right answer from the matrix I obtain by taking A - 2(lambda)*I

OpenStudy (anonymous):

send me your e-mail, or please write to: diegoesdiego@hotmail.com. I will return the complete procedure. bye.

OpenStudy (amistre64):

\[\left[\begin{matrix}1-2 & 0 & 0&0 \\ 1&2-2 &0&0 \\ 1&2&-1-2&1 \\ 1&1&0&4-2\end{matrix}\right]\] \[\left[\begin{matrix}-1 & 0 & 0&0 \\ 1&0 &0&0 \\ 1&2&-3&1 \\ 1&1&0&2\end{matrix}\right]\] and row reduce

OpenStudy (anonymous):

what is row reduce though?

OpenStudy (amistre64):

getting it into echelon form ... how did you do the others?

OpenStudy (anonymous):

set up systems of equations by multiplying with variables for the other part

OpenStudy (amistre64):

it row reduces to {1, 0, 0, 0} {0, 1, 0, 2} {0, 0, 1, 1} {0, 0, 0, 0} and negate the last column since there is only 1 free variable

OpenStudy (amistre64):

well, the last one will not be a zero, itll be a 1

OpenStudy (amistre64):

x = w{ 0} y = w{-2} z = w{-1} w = w{ 1}

OpenStudy (anonymous):

I'll trust you on that one. I'll have to look into that some more. honestly, I've never seen that. That does make it come out right. Thanks

OpenStudy (amistre64):

{-1, 0, 0, 0} *-1 { 1, 0, 0, 0} { 1, 2, -3, 1} { 1, 1, 0, 2} { 1, 0, 0, 0} { 1, 0, 0, 0} add the rows { 1, 2, -3, 1} using the top row { 1, 1, 0, 2} { 1, 0, 0, 0} { 0, 0, 0, 0} rearrange in a more echelon form { 0, 2, -3, 1} { 0, 1, 0, 2} { 1, 0, 0, 0} { 0, 1, 0, 2} *-2 and add to the 3rd row { 0, 2, -3, 1} { 0, 0, 0, 0} { 1, 0, 0, 0} { 0, 1, 0, 2} { 0, 0, -3, -3} /-3 { 0, 0, 0, 0} { 1, 0, 0, 0} { 0, 1, 0, 2} finished results { 0, 0, 1, 1} { 0, 0, 0, 0} ^^ free variable

OpenStudy (amistre64):

define the rest of them in terms of the free variable. which i did above